# Knowledge Entropy Decay during Language Model Pretraining Hinders New Knowledge Acquisition

### v1

@@ Abstract
In this work, we investigate how a model’s tendency to broadly integrate its parametric knowledge evolves throughout pretraining, and how this behavior affects overall performance, particularly in terms of knowledge acquisition and forgetting. We introduce the concept of knowledge entropy, which quantifies the range of memory sources the model engages with; high knowledge entropy indicates that the model utilizes a wide range of memory sources, while low knowledge entropy suggests reliance on specific sources with greater certainty. Our analysis reveals a consistent decline in knowledge entropy as pretraining advances. We also find that the decline is closely associated with a reduction in the model’s ability to acquire and retain knowledge, leading us to conclude that diminishing knowledge entropy (smaller number of active memory sources) impairs the model’s knowledge acquisition and retention capabilities. We find further support for this by demonstrating that increasing the activity of inactive memory sources enhances the model’s capacity for knowledge acquisition and retention [^1] Code in https://github.com/kaistAI/Knowledge-Entropy.git.
@@ 1 Introduction
Recent works have analyzed how language models store world knowledge in their parameters and utilize this knowledge to generate responses during inference time (Geva et al., 2021; Dai et al., 2022a; b; Meng et al., 2022; Yao et al., 2024). However, little is known about how their behavior of integrating various factual knowledge embedded in their parameters changes throughout the pretraining stage. In this work, we perform deep analysis into how a model’s property of broadly integrating diverse parametric knowledge evolves throughout pretraining and how these shifts affect overall performance, particularly in terms of knowledge acquisition and forgetting in a continual learning setup. We hypothesize that this varying level of integration may explain why models in the later stage of pretraining encounter challenges in acquiring new knowledge (Dohare et al., 2024; Jang et al., 2022; Chang et al., 2024).
@@
We introduce knowledge entropy, which reflects how a language model integrates various knowledge sources, to investigate how this behavior evolves throughout pretraining. Recent studies have shown that feed-forward layers (FFNs) serve as a key-value memory (Geva et al., 2022; 2021; Dai et al., 2022a). Building on this research, as shown in Figure 1, we view the second projection matrix $V$ as a memory, composed of memory vectors, which store the model’s parametric knowledge, and view the first projection matrix $K$ as generating coefficients $\bar{C}$ that determine how these memory vectors are combined. Knowledge entropy measures how sparsely these memory coefficients are distributed; high knowledge entropy indicates that the model tends to integrate broad memory vectors whereas the model with low knowledge entropy uses certain memory vector with high certainty. We analyze models at different stages of pretraining to investigate how knowledge entropy changes throughout pretraining. Our findings show that models in the later stages of pretraining tend to exhibit lower knowledge entropy, suggesting a shift from utilizing a larger set of active memory vectors to a smaller, more focused set as pretraining progresses.
@@
We hypothesize that knowledge entropy change would influences the model’s behavior when encountering new knowledge. To test this, we conduct a thorough analysis of the model’s ability to acquire new knowledge and retain existing knowledge in a continual knowledge learning [^2] In this work, we use the term “continual knowledge learning” and “continual learning” interchangeably. scenario on a target corpora (Jang et al., 2022; Wu et al., 2023), starting from different stages of pretraining. This involves further training the pretrained model on new-domain corpora using a language modeling objective to integrate new knowledge. Results show a strong correlation between knowledge entropy and the model’s ability to acquire and retain knowledge: both knowledge entropy and knowledge acquisition and retention decrease as the pretraining progresses.
@@
We assume that this correlation arises because lower knowledge entropy indicates a smaller set of active memory vectors, which leads to frequent overwriting on these memory vectors to store new knowledge. To test this assumption, we conduct experiments where we artificially increase the activity of previously inactive memory vectors, allowing the model to store new knowledge across a broader range of memory vectors. Surprisingly, we observe that these modified models demonstrate improved knowledge acquisition and reduced forgetting compared to the original models when undergo continual knowledge learning; though not to the same extent as in the original pretrained model with equivalent knowledge entropy. Such a result bolsters our hypothesis that having a limited number of active memory vectors (low knowledge entropy) plays a critical part in explaining the degradation of the model’s ability to acquire and retain knowledge as pretraining advances.
@@
Overall, our findings reveal that as pretraining progresses, models exhibit narrower integration of memory vectors, reflected by decreasing knowledge entropy, which hinders both knowledge acquisition and retention. Models in the later stages of pretraining [^3] We define the final stage as the last stage of the pre-determined learning rate schedule, with the most decayed learning rate. The mid-stage refers to around 50% of the schedule, while the initial stage refers to around 20%., show low knowledge entropy, leading to poor knowledge acquisition and higher forgetting rates despite being trained on larger datasets. In contrast, early-stage models display high knowledge entropy, allowing them better in knowledge acquisition and retention; however, their performance is often limited by weaker language modeling capabilities. Thus, mid-stage models strike a balance, showing strong knowledge acquisition and retention along with overall performance, making them a practical choice for further training to inject new knowledge. To the best of our knowledge, this is the first work to analyze how a model’s behavior in integrating various memory vectors changes across pretraining stages and the subsequent effects on performance when acquiring new knowledge in a continual knowledge learning setup.
2 Related Works
@@
Recent studies have shown that language models embed world knowledge within their parameters and integrate this knowledge to generate responses (Yang, 2024; Petroni et al., 2019; Wang et al., 2021). Thereby, various research aims to understand these dynamics of knowledge in language models (how they learn, store, and engage their parametric knowledge) during inference and training phases. Several works have focused on investigating the inference process: Geva et al. (2023) analyzes the role of different layers in language models. Allen-Zhu & Li (2024b) demonstrates that model parameters have a limited knowledge capacity. Some studies suggest key-value memory (Geva et al., 2021; Meng et al., 2022; Dai et al., 2022a). Other research focuses on the pretraining phase. Chang et al. (2024) explores how language models acquire knowledge. Liu et al. (2021) studies the sequence that language models learn various types of knowledge. Allen-Zhu & Li (2024a) examines strategies to enhance knowledge storage and extraction. Teehan et al. analyzes internal structural changes. Sun & Dredze (2024) investigates the interaction between pretraining and finetuning. Our work is the first to explore how the behavior of language models in integrating their knowledge evolves throughout the pretraining phase, and to analyze how these changes affect model performance in terms of knowledge acquisition and forgetting in continual knowledge learning.
@@
In information theory, entropy quantifies the value of information, where predictable (certain) events have low entropy and unpredictable (uncertain) events have high entropy (Lairez, 2022; Majenz, 2018). In natural language processing (NLP), entropy is used in various ways to measure the certainty of language models. Yang (2024) analyzes the entropy of model outputs based on input prompts. Araujo et al. (2022) calculates the entropy of outputs at each layer to determine weight adjustments in a continual learning setup. Other works focus on token probability entropy to understand the information required to predict the next word in a sequence (Vazhentsev et al., ; Geng et al., 2024; Malinin & Gales, 2021). Lower entropy in a model’s predictions may indicate that the model has become more certain about its predictions based on training data. Additionally, Kumar & Sarawagi (2019) measures entropy over the cross-attention layer to assess the uncertainty in the attention layer of encoder-decoder models. The entropy proposed in our paper, knowledge entropy, differs from previous works in that it focuses on the entropy of a model’s parametric knowledge, assessing the uncertainty or variability in utilizing the knowledge encoded within the language model.
@@ 3 Knowledge Entropy
In this section, we introduce knowledge entropy (Section 3.1) to examine how broadly the model integrates its parametric knowledge and an experiment setup to measure it (Section 3.2). Next, in Section 3.3, we measure knowledge entropy at various pretraining stages to analyze how the model’s knowledge integration behavior evolves over pretraining. In Section 3.4, we extend our investigation by exploring alternative definitions of entropy.
@@ 3.1 Definition
In this work, we introduce a new concept, knowledge entropy, to analyze the scope of a model’s access patterns to its parametric knowledge. Low knowledge entropy suggests that the model relies more on a narrower set of specific knowledge sources with high certainty whereas high knowledge entropy indicates that the model integrates with a diverse range of knowledge sources. Inspired by prior works that consider feed-forward layers (FFNs) as key-value memory containing model’s parametric memory (Geva et al., 2021; Dai et al., 2022a; Meng et al., 2022; Dong et al., 2022), we consider the knowledge source to be the memory vectors, which is the second projection matrix of FFN, and measurement of how broadly the model integrates as memory coefficients, calculated by the first projection matrix.
@@
$C^{(l)}_{n,j}$ represents the coefficient of $j$-th token position of $n$-th instance at layer $l$, $\bar{c}^{l}_{i}$ indicates an $i$-th element from $\bar{C}^{l}$, $T_{n}$ is the sequence length of the $n$-th instance in the dataset $\mathcal{D}$, $m$ is the inner dimension of feed-forward layer, and $L$ denotes the number of layers in the model.
@@ 3.2 Experiment Setup
In the case of OLMo, the memory coefficient $C^{(l)}_{n,j}$ is calculated as $C^{(l)}_{n,j}=\text{abs}(\text{SwiGLU}(\mathbf{x_{j}}))$ where $\mathbf{x_{j}}$ is the j-th token of input $\mathbf{x}$ and SwiGLU (Shazeer, 2020) is the activation function. We apply the absolute value since the SwiGLU allows negative values and the magnitude determines the contribution of the corresponding memory vector in the linear combination. Then, the absolute values are converted into probability distribution. We also experiment that the trend persists with different choices of activation functions. Further details regarding knowledge entropy can be found in Appendix A.1.
@@ 3.3 Final models tend to exhibit lower knowledge entropy
Figure 3 illustrates how knowledge entropy (y-axis) changes across different stages of pretraining (x-axis). The results show a consistent decrease in knowledge entropy as pretraining progresses in both 1B and 7B models. This trend suggests that models in the later stage of pretraining tends to engage with a narrower range of memories, relying more heavily on specific memory vectors rather than accessing and integrating knowledge from a broader range of memories. Consistent reduction in knowledge entropy is observed across all layers, with the most significant reduction occuring in the last layer, which closely resembles the output distribution right before the token prediction (Figure  8 in Appendix).
@@ 3.4 Similar Trends are observed by different definitions of Entropy
While our work defines knowledge entropy focusing on the feed-forward layer, previous studies have examined entropy in different contexts, such as the entropy of attention (Kumar & Sarawagi, 2019) and the entropy of next token prediction over the vocabulary space (Vazhentsev et al., ; Malinin & Gales, 2021). To gain a more comprehensive understanding of the model’s overall behavior, we extend our analysis by exploring the entropy trends in both attention mechanisms and next token prediction. The formula and details are in Appendix A.2)
@@
Following Kumar & Sarawagi (2019), we measure attention entropy to capture the degree of uncertainty in attention weight. It is calculated as the sum of layer-wise entropy, where the layer-wise entropy measures the sparsity of the attention weight in each attention head. Thus, attention entropy reflects how much weight the model assigns to specific tokens with confidence when generating the next token given the input based on token relationships. Figure 3 shows that attention entropy consistently decreases during pretraining, with a sharp decline in the early stages followed by a more gradual reduction. This trend suggests that the model learns to focus on contextually important tokens within the attention layer.
@@
Entropy can also be measured based on the probability distribution of the next token prediction over the vocabulary space (Vazhentsev et al., ; Geng et al., 2024; Malinin & Gales, 2021). Figure 3 shows that the entropy of the next token prediction also consistently decreases throughout pretraining, reflecting the model’s increasing certainty in its next token prediction.
@@ 4 Knowledge Acquisition and Forgetting
We hypothesize that the reduction of knowledge entropy as pretraining progresses impacts the model’s knowledge acquisition and forgetting as low knowledge entropy indicates sparse activation of memory vectors, thus the vectors are likely to be consistently overwritten when new knowledge is introduced. To test this hypothesis, we measure knowledge acquisition and forgetting using checkpoints from different stages of pretraining in a continual knowledge learning setup (Jang et al., 2022; Wu et al., 2023), where further training is performed on new-domain corpora by next token prediction to inject new knowledge to the pretrained models. Section 4.1 details the experimental setup and the metrics used. In Section 4.2, we present the results of knowledge acquisition and forgetting across various pretraining stages. Section 4.3 further explores whether a relationship exists between the two behaviors: activating the inactive memory vectors increases the knowledge acquisition ability.
@@ 4.1 Experiment Setup
We experiment using intermediate checkpoints from OLMo [^6] We select these intermediate checkpoints following Sun & Dredze (2024), which explores language model (OLMo) throughout pretraining. Available checkpoints can be found here.. Hyperparameters are chosen following previous research on continual knowledge learning (Jang et al., 2022; Kim et al., 2023) and we test various combinations to assess their generalizability. For batch size, we test 128 and 2048; for learning rate, we experiment with 1e-4, 4e-4, and 1e-3. We also investigate the effect of training duration by comparing a single epoch to three epochs. Among these configurations, we focus primarily on batch size 128, learning rate 4e-4, and single epoch training as this setup most closely aligns with continual knowledge learning.
@@
We experiment on a subset of two datasets [^7] We randomly sample 205k instances for each dataset.: PubMed  [^8] Datasets in huggingface, a corpus of bio-medical and life science topics with abstracts, and C4 (Raffel et al., 2020), a large-scale corpus comprising diverse text data gathered from web pages. We use PubMed as the primary dataset as it contains more new knowledge, making it a better fit for our continual knowledge learning setup (Appendix B.1). In addition to the dataset, we inject synthetic knowledge during training to assess the model’s ability to acquire new information. Specifically, we inject Fictional Knowledge dataset (Chang et al., 2024), which is designed to assess how well language models acquire factual knowledge during pretraining [^9] We slightly modified the dataset to our setting of which details are in Appendix B.1. This dataset includes 130 paragraphs about fictional yet realistic entities and 1,950 probes where each paragraphs contain 15 different probes. The passage is incorporated into the training batch 10 times during the continual knowledge learning. After training, we evaluate the models on evaluation probes of the Fictional Knowledge dataset to measure knowledge acquisition, and evaluate on six downstream tasks in zero-shot manner (Sun & Dredze, 2024; Groeneveld et al., 2024) to measure knowledge forgetting (SciQ (Welbl et al., 2017), Winogrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), OBQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), and ARC Easy (Clark et al., 2018)). Detailed explanation is included in Appendix B.1.
@@
Knowledge acquisition of a language model $\theta$ is measured with the probing performance on evaluation probes following Chang et al. (2024). When given the injected knowledge $C$, each instance $c_{i}$ in a corpus has a corresponding set of probes $P_{c_{i}}$, containing 15 different probes. To measure how well the model recalls the injected knowledge, we compute the probe performance $\mathcal{K}(\theta)$, the average log probability $\ell$($p_{i}$;$\theta$) of target span for each probe $p_{i}\in P_{c_{i}}$ across all instances in $c_{i}\in C$ and calculate average; $\mathcal{K}(\theta)=\frac{1}{|C|}\sum_{c_{i}\in C}\frac{1}{|P_{c_{i}}|}\sum_{p_{i}\in P_{c_{i}}}\ell(p_{i};\theta)$. The knowledge acquisition metric $\mathcal{A}(\theta)$ is defined as the improvement rate of $\mathcal{K}(\theta)$ from $\theta_{\text{PT}}$ to $\theta_{\text{CL}}$, where $\theta_{\text{PT}}$ represents the model checkpoint from a pretraining step, which serves as the starting point and $\theta_{\text{CL}}$ represents the model after continual knowledge learning. High $\mathcal{A}(\theta)$ indicates the model has learned new knowledge well.
To measure knowledge forgetting of a language model, we measure average performance over six downstream tasks $\mathcal{P}(\theta)$. Knowledge forgetting $\mathcal{F}(\theta)$ is calculated by the reduction rate from $\theta_{\text{PT}}$ to $\theta_{\text{CL}}$. Low $\mathcal{F}(\theta)$ indicates the models have retained its existing knowledge. Equation and detailed explanation is presented in Appendix B.2.

### v3

@@ Abstract
In this work, we investigate how a model’s tendency to broadly integrate its parametric knowledge evolves throughout pretraining and how this behavior affects overall performance, particularly in terms of knowledge acquisition and forgetting. We introduce the concept of knowledge entropy, which quantifies the range of memory sources the model engages with; high knowledge entropy indicates that the model utilizes a wide range of memory sources, while low knowledge entropy suggests reliance on specific sources with greater certainty. Our analysis reveals a consistent decline in knowledge entropy as pretraining advances. We also find that the decline is closely associated with a reduction in the model’s ability to acquire and retain knowledge, leading us to conclude that diminishing knowledge entropy (smaller number of active memory sources) impairs the model’s knowledge acquisition and retention capabilities. We find further support for this by demonstrating that increasing the activity of inactive memory sources enhances the model’s capacity for knowledge acquisition and retention. [^1] Code in https://github.com/kaistAI/Knowledge-Entropy.git
@@ 1 Introduction
Recent studies have analyzed how language models store world knowledge in their parameters and utilize this knowledge to generate responses during inference time (Geva et al., 2021; Dai et al., 2022a; b; Meng et al., 2022; Yao et al., 2024). However, little is known about how their behavior of integrating various factual knowledge embedded in their parameters changes throughout the pretraining stage. In this work, we perform a deep analysis of how a model’s property of broadly integrating diverse parametric knowledge evolves throughout pretraining and how these shifts affect overall performance, particularly in terms of knowledge acquisition and forgetting in a continual learning setup. We hypothesize that this varying level of integration may explain why models in the later stages of pretraining encounter challenges in acquiring new knowledge (Dohare et al., 2024; Jang et al., 2022; Chang et al., 2024).
We introduce knowledge entropy, which reflects how a language model integrates various knowledge sources, to investigate how this behavior evolves throughout pretraining. Recent studies have shown that feed-forward layers (FFNs) serve as a key-value memory (Geva et al., 2022; 2021; Dai et al., 2022a). Building on this research, as shown in Figure 1, we view the second projection matrix $V$ as a memory, composed of memory vectors, which store the model’s parametric knowledge, and view the first projection matrix $K$ as generating coefficients $\bar{C}$ that determine how these memory vectors are combined. Knowledge entropy measures how sparsely these memory coefficients are distributed; high knowledge entropy indicates that the model tends to integrate a broad range of memory vectors whereas the model with low knowledge entropy relies on specific memory vectors with high certainty. We analyze models at different stages of pretraining to investigate how knowledge entropy changes throughout pretraining. Our findings show that models in the later stages of pretraining tend to exhibit lower knowledge entropy, suggesting a shift from utilizing a larger set of active memory vectors to a smaller, more focused set as pretraining progresses.
@@
We hypothesize that changes in knowledge entropy would influence the model’s behavior when encountering new knowledge. To test this, we conduct a thorough analysis of the model’s ability to acquire new knowledge and retain existing knowledge in a continual knowledge learning [^2] In this work, we use the term “continual knowledge learning” and “continual learning” interchangeably. scenario on a target corpus (Jang et al., 2022; Wu et al., 2023), starting from different stages of pretraining. This involves further training the pretrained model on new-domain corpora using a language modeling objective to integrate new knowledge. Our results reveal a strong correlation between knowledge entropy and the model’s ability to acquire and retain knowledge: both knowledge entropy and knowledge acquisition and retention decrease as the pretraining progresses.
@@
We assume that this correlation arises because lower knowledge entropy indicates a smaller set of active memory vectors, leading to frequent overwriting of these memory vectors to store new knowledge. To test this assumption, we conduct experiments where we artificially increase the activity of previously inactive memory vectors, allowing the model to store new knowledge across a broader range of memory vectors. Surprisingly, we observe that these modified models demonstrate improved knowledge acquisition and reduced forgetting compared to the original models when undergoing continual knowledge learning; though not to the same extent as the original pretrained model with equivalent knowledge entropy. Such a result bolsters our hypothesis that having a limited number of active memory vectors (low knowledge entropy) plays a critical role in explaining the degradation of the model’s ability to acquire and retain knowledge as pretraining advances.
@@
Overall, our findings reveal that as pretraining progresses, models exhibit a narrower integration of memory vectors, reflected by decreasing knowledge entropy, which hinders both knowledge acquisition and retention. Models in the later stages of pretraining [^3] We define the final stage as the last stage of the pre-determined learning rate schedule, with the most decayed learning rate. The mid-stage refers to around 50% of the schedule, while the initial stage refers to around 20%. show low knowledge entropy, leading to poor knowledge acquisition and higher forgetting rates despite being trained on larger datasets. In contrast, early-stage models display high knowledge entropy, enabling better knowledge acquisition and retention; however, their performance is often limited by weaker language modeling capabilities. Thus, mid-stage models strike a balance, showing strong knowledge acquisition and retention along with overall performance, making them a practical choice for further training to incorporate new knowledge. To the best of our knowledge, this is the first work to analyze how a model’s behavior in integrating various memory vectors changes across pretraining stages and the subsequent effects on performance when acquiring new knowledge in a continual knowledge learning setup.
2 Related Work
@@
Recent studies have shown that language models embed world knowledge within their parameters and integrate this knowledge to generate responses (Yang, 2024; Petroni et al., 2019; Wang et al., 2021). Thereby, various research efforts aim to understand these dynamics of knowledge in language models (how they learn, store, and engage their parametric knowledge) during inference and training phases. Several studies have focused on investigating the inference process: Geva et al. (2023) analyzes the role of different layers in language models. Allen-Zhu & Li (2024b) demonstrates that model parameters have a limited knowledge capacity. Some studies suggest key-value memory (Geva et al., 2021; Meng et al., 2022; Dai et al., 2022a). Other research focuses on the pretraining phase. Liu et al. (2021) studies the sequence that language models learn various types of knowledge. Allen-Zhu & Li (2024a) examines strategies to enhance knowledge storage and extraction. Teehan et al. analyzes internal structural changes. Sun & Dredze (2024) investigates the interaction between pretraining and finetuning. Chang et al. (2024) analyzes patterns of knowledge acquisition specifically during the pretraining process, addressing the question of how language models acquire knowledge during pretraining. While their study shares similarities with ours in investigating knowledge acquisition behavior during LLM training, our work takes a different focus. We aim to understand why LLMs encounter increasing difficulty in acquiring new knowledge as pretraining progresses, exploring the underlying reasons behind the challenges faced by later-stage models in learning new knowledge. To the best of our knowledge, our work is the first to explore how the behavior of language models in integrating their knowledge evolves throughout the pretraining phase, and to analyze how these changes affect model performance in terms of knowledge acquisition and forgetting in continual knowledge learning.
@@
In information theory, entropy quantifies the value of information, where predictable (certain) events have low entropy and unpredictable (uncertain) events have high entropy (Lairez, 2022; Majenz, 2018). In natural language processing (NLP), entropy is used in various ways to measure the certainty of language models. Yang (2024) analyzes the entropy of model outputs based on input prompts. Araujo et al. (2022) calculates the entropy of outputs at each layer to determine weight adjustments in a continual learning setup. Other papers focus on token probability entropy to understand the information required to predict the next word in a sequence (Vazhentsev et al., ; Geng et al., 2024; Malinin & Gales, 2021). Lower entropy in a model’s predictions may indicate that the model has become more certain about its predictions based on training data. Additionally, Kumar & Sarawagi (2019) measures entropy over the cross-attention layer to assess the uncertainty in the attention layer of encoder-decoder models. The entropy proposed in our paper, knowledge entropy, differs from previous work in that it focuses on the entropy of a model’s parametric knowledge, assessing the uncertainty or variability in utilizing the knowledge encoded within the language model.
@@ 3 Knowledge Entropy
In this section, we introduce knowledge entropy (Section 3.1) to examine how broadly the model integrates its parametric knowledge and describe the experimental setup used to measure it (Section 3.2). Next, in Section 3.3, we measure knowledge entropy at various pretraining stages to analyze how the model’s knowledge integration behavior evolves over pretraining. In Section 3.4, we extend our investigation by exploring alternative definitions of entropy.
@@ 3.1 Definition
In this work, we introduce a new concept, knowledge entropy, to analyze the scope of a model’s access patterns to its parametric knowledge. Low knowledge entropy suggests that the model relies on a narrower set of specific knowledge sources with high certainty whereas high knowledge entropy indicates that the model integrates with a diverse range of knowledge sources. Inspired by prior research that considers feed-forward layers (FFNs) as key-value memory containing a model’s parametric memory (Geva et al., 2021; Dai et al., 2022a; Meng et al., 2022; Dong et al., 2022), we consider the knowledge source to be the memory vectors, which is the second projection matrix of FFN. We measure how broadly the model integrates these memory vectors with memory coefficients, which are calculated by the first projection matrix and the activation function.
@@
$C^{(l)}_{n,j}$ represents the coefficient of the $j$-th token position of the $n$-th instance at layer $l$, $\bar{c}^{l}_{i}$ indicates the $i$-th element from $\bar{C}^{l}$, $T_{n}$ is the sequence length of the $n$-th instance in the dataset $\mathcal{D}$, $m$ is the inner dimension of feed-forward layer, and $L$ denotes the number of layers in the model.
@@ 3.2 Experiment Setup
In the case of OLMo, the memory coefficient $C^{(l)}_{n,j}$ is calculated as $C^{(l)}_{n,j}=\text{abs}(\text{SwiGLU}(\mathbf{x_{j}}))$ where $\mathbf{x_{j}}$ is the j-th token of input $\mathbf{x}$ and SwiGLU (Shazeer, 2020) is the activation function. We apply the absolute value since the SwiGLU allows negative values and the magnitude determines the contribution of the corresponding memory vector in the linear combination. Then, the absolute values are converted into a probability distribution. We also show experimentally that the trend persists with different choices of activation functions. Further details regarding knowledge entropy can be found in Appendix A.2.
@@ 3.3 Final Models Tend to Exhibit Lower Knowledge Entropy
Figure 3 illustrates how knowledge entropy (y-axis) changes across different stages of pretraining (x-axis). The results show a consistent decrease in knowledge entropy as pretraining progresses in both 1B and 7B models. This trend suggests that models in the later stages of pretraining tend to engage with a narrower range of memories, relying more heavily on specific memory vectors rather than accessing and integrating knowledge from a broader range of memories. A consistent reduction in knowledge entropy is observed across all layers, with the most significant reduction occurring in the last layer, which closely resembles the output distribution right before the token prediction (Figure  8 in Appendix).
@@ 3.4 Similar Trends are Observed by Different Definitions of Entropy
While our work defines knowledge entropy focusing on the feed-forward layer, previous studies have examined entropy in different contexts, such as the entropy of attention (Kumar & Sarawagi, 2019) and the entropy of next-token prediction over the vocabulary space (Vazhentsev et al., ; Malinin & Gales, 2021). To gain a more comprehensive understanding of the model’s overall behavior, we extend our analysis by exploring the entropy trends in both attention mechanisms and next-token prediction. The formula and details are provided in Appendix A.3)
@@
Following Kumar & Sarawagi (2019), we measure attention entropy to capture the degree of uncertainty in attention weights. It is calculated as the sum of layer-wise entropy, where the layer-wise entropy measures the sparsity of the attention weights in each attention head. Thus, attention entropy reflects how much weight the model assigns to specific tokens with confidence when generating the next token given the input based on token relationships. Figure 3 shows that attention entropy consistently decreases during pretraining, with a sharp decline in the early stages followed by a more gradual reduction. This trend suggests that the model learns to focus on contextually important tokens within the attention layer.
@@
Entropy can also be measured based on the probability distribution of next-token predictions over the vocabulary space (Vazhentsev et al., ; Geng et al., 2024; Malinin & Gales, 2021). Figure 3 shows that the entropy of the next token prediction also consistently decreases throughout pretraining, reflecting the model’s increasing certainty in its next token prediction.
@@ 4 Knowledge Acquisition and Forgetting
We hypothesize that the reduction of knowledge entropy as pretraining progresses impacts the model’s knowledge acquisition and forgetting as low knowledge entropy indicates sparse activation of memory vectors, thus the vectors are likely to be consistently overwritten when new knowledge is introduced. To test this hypothesis, we measure knowledge acquisition and forgetting using checkpoints from different stages of pretraining in a continual knowledge learning setup (Jang et al., 2022; Wu et al., 2023), where further training is performed on new-domain corpora by next token prediction to inject new knowledge into the pretrained models. Section 4.1 details the experimental setup and the metrics used. In Section 4.2, we present the results of knowledge acquisition and forgetting across various pretraining stages. Finally, Section 4.3 further explores whether a relationship exists between the two behaviors: activating the inactive memory vectors increases the knowledge acquisition ability.
@@ 4.1 Experiment Setup
We experiment using intermediate checkpoints from OLMo [^6] We select these intermediate checkpoints following Sun & Dredze (2024), which explores language model (OLMo) throughout pretraining. Available checkpoints can be found here.. Hyperparameters are chosen following previous research on continual knowledge learning (Jang et al., 2022; Kim et al., 2023) and we test various combinations to assess their generalizability. For batch size, we test 128 and 2048; for learning rate, we experiment with 1e-4, 4e-4, and 1e-3. We also investigate the effect of training duration by comparing a single epoch to three epochs. Among these configurations, we focus primarily on a batch size of 128, a learning rate of 4e-4, and single-epoch training as this setup most closely aligns with continual knowledge learning.
@@
We experiment on a subset of two datasets [^7] We randomly sample 205k instances for each dataset.: PubMed  [^8] Datasets in huggingface, a corpus of bio-medical and life science topics with abstracts, and C4 (Raffel et al., 2020), a large-scale corpus comprising diverse text data gathered from web pages. We use PubMed as the primary dataset as it contains more new knowledge, making it a better fit for our continual knowledge learning setup (Appendix B.1). In addition to the dataset, we inject synthetic knowledge during training to assess the model’s ability to acquire new information. Specifically, we utilize Fictional Knowledge dataset (Chang et al., 2024), which is designed to assess how well language models acquire factual knowledge during pretraining [^9] We slightly modified the dataset to our setting of which details are in Appendix B.1. This dataset includes 130 paragraphs about fictional yet realistic entities and 1,950 probes where each paragraph contains 15 different probes. The passages are incorporated into the training batch 10 times during the continual knowledge learning. After training, we evaluate the models on evaluation probes of the Fictional Knowledge dataset to measure knowledge acquisition, and evaluate on six downstream tasks in zero-shot manner (Sun & Dredze, 2024; Groeneveld et al., 2024) to measure knowledge forgetting (SciQ (Welbl et al., 2017), Winogrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), OBQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), and ARC Easy (Clark et al., 2018)). Detailed explanation is included in Appendix B.1.
@@
Knowledge acquisition of a language model $\theta$ is measured with the probing performance on evaluation probes following Chang et al. (2024). When given the injected knowledge $W$, each instance $w_{i}$ in a corpus has a corresponding set of probes $P_{w_{i}}$, containing 15 different probes. To measure how well the model recalls the injected knowledge, we compute the probe performance $\mathcal{K}(\theta)$, the average log probability $\ell$($p_{i}$;$\theta$) of the target span for each probe $p_{i}\in P_{w_{i}}$ across all instances in $w_{i}\in W$ and calculate average; $\mathcal{K}(\theta)=\frac{1}{|W|}\sum_{w_{i}\in W}\frac{1}{|P_{w_{i}}|}\sum_{p_{i}\in P_{w_{i}}}\ell(p_{i};\theta)$. The knowledge acquisition metric $\mathcal{A}(\theta)$ is defined as the improvement rate of $\mathcal{K}(\theta)$ from $\theta_{\text{PT}}$ to $\theta_{\text{CL}}$, where $\theta_{\text{PT}}$ represents the model checkpoint from a pretraining step, which serves as the starting point and $\theta_{\text{CL}}$ represents the model after continual knowledge learning. High $\mathcal{A}(\theta)$ indicates the model has learned new knowledge well.
To measure knowledge forgetting of a language model, we measure average performance over six downstream tasks $\mathcal{P}(\theta)$. Knowledge forgetting $\mathcal{F}(\theta)$ is calculated by the reduction rate from $\theta_{\text{PT}}$ to $\theta_{\text{CL}}$. Low $\mathcal{F}(\theta)$ indicates that the models have retained their existing knowledge. The equation and a detailed explanation are presented in Appendix B.2.